OpenStudy (anonymous):

Sketch the region enclosed by the lines: x = 0, x = 6, y = 2, and y = 6. 6. Find the area of the region.

OpenStudy (mathmale):

Hints: Equations such as x = 0 and x = 6 represent vertical lines, and those such as y = 2 and y = 6 represent horizontal lines. Hope that helps you get started graphing this situation. a graph will help you visualize the length and width of this region. Later you'll be able to read off the length and the width of the region almost immediately, without a drawing or any but the simplest calculations.

OpenStudy (anonymous):

How would I use those to graph the situation? If I knew how the get the lengths, I could solve the problem, no problem.

OpenStudy (anonymous):

the lengths are the difference 6-0=6 and 6-2=4

OpenStudy (anonymous):

So to find the area, 6*4=24?

OpenStudy (anonymous):

\[x_2-x_1\] \[y_2-y_1\]

OpenStudy (anonymous):

yes,but did you draw the shape

OpenStudy (anonymous):

24 is correct

OpenStudy (anonymous):

No, but I assume it's a square or rectangle.

OpenStudy (anonymous):

rectangle

OpenStudy (anonymous):

Oh yeah, the lengths are not equal. duh moment. oops.

OpenStudy (anonymous):



OpenStudy (mathmale):

Very nice graph, @jonask! OrangeMaster: Just by "eye-balling" this sketch, you can probably tell that the length of this rectangle is 6 (that is, 6-0=6). What is the width of this rectangle? Consider this to be just a summary, because you've already indicated the area correctly: 4*6 = 24 square units.